Expand as much as possible log(10,000x).
Answer by AnlytcPhil(1116)

Expand as much as possible log(10,000x).
The laws of logs are
(1) log(uv) = log(u) + log(v)
(2) log(u/v) = log(u) - log(v)
(3) log(uv) = v·log(v)
(4) log(10) = 1
log(10000x) equals
log(104x) because 10000 = 104, which equals
log(104) + log(x) by (1) above, which equals
4·log(10) + log(x) by (3) above, which equals
4·1 + log(x) by (4) above, which equals
4 + log(x)
Edwin